# Consolidating and Developing Benchmarking Datasets for the Nepali Natural Language Understanding Tasks

### v2

@@ Abstract
Jinu Nyachhyon, Mridul Sharma, Prajwal Thapa, Bal Krishna Bal  Information and Language Processing Research Lab (ILPRL), Kathmandu University Corresponding Author: nyachhyonjinu@gmail.com
@@ 1 Introduction
SA 65.1K 16.3K Sentiment Macro F1, Acc Movie Reviews
@@
Nepali is written in the Devanagari script and is a highly inflected language.The Nepali language incorporates a complex system of noun, adjective, and verb inflections, including gender, case, and number Bal (2004). It has a rich vocabulary with many homonyms and is spoken in different dialects across various regions, and there are variations in vocabulary, grammar, and pronunciation. Developing and establishing robust models for Nepali requires reliable methods to evaluate their quality and effectiveness and it is essential to have tools that can assess how well these models address the language’s unique linguistic challenges while identifying their limitations.
@@ 2 Related Works
Nepali Sentiment Analysis (NepSA) Singh et al. (2020) is a targeted aspect-based sentiment analysis dataset, comprising 3,068 comments extracted from 37 YouTube videos across 9 channels. The dataset is annotated using a binary sentiment polarity schema across six aspects: General, Profanity, Violence, Feedback, Sarcasm, and Out-of-scope. Another dataset, Aspect-Based Sentiment Analysis Tamrakar et al. (2020), contains 1,576 sentences, equally divided between positive and negative sentiments. Additionally, sentiment analysis datasets like Nepali Language Sentiment Analysis - Movie Reviews Ghimire with 602 data points, and Nepali Sentiment Analysis Acharya with 2,161 data points found on Kaggle, are limited in size and domain-specific. For our benchmark, we utilized the NepCOV19Tweets dataset Sitaula et al. (2021), which includes ~33.5k sentiments labeled as positive, negative, or neutral. From this, we selected 14.9k positive and 13.5k negative data points for the SA dataset. A more recent dataset, Sentiment of Election-Based Nepali Tweets Pokharel , contains ~17.8k tweets but includes English characters and numbers, making it less suitable for our benchmarked dataset. To our knowledge, there are no publicly available datasets for coreference resolution, acceptability judgment, paraphrase and similarity detection, commonsense reasoning, pronoun coreference resolution, general masked evaluation, or NLI in the Nepali language. Despite some studies focusing on Nepali grammar, the lack of datasets for these advanced tasks limits the development of comprehensive NLU benchmarks.
@@ 8 Limitations
First, the datasets introduced in the NLUE benchmark were primarily created by translating existing English-language datasets from benchmarks such as GLUE and xGLUE, using automated tools like GPT-4o-mini and Gemini2.5-flash, supplemented by manual corrections. Although efforts were made to ensure translation accuracy, subtle linguistic nuances, cultural contexts, and idiomatic expressions specific to Nepali may not have been fully captured. The small size of certain datasets (e.g., CR and RTE) limits model performance and shows models’ lack of generalization in smaller datasets. Second, the study evaluates a range of models with varying parameter sizes (67M to 276M), but resource constraints prevented the inclusion of larger, state-of-the-art models or extensive hyperparameter tuning. Finally, the reliance on specific evaluation metrics (e.g., accuracy, Macro-F1 score, Spearman, and Pearson correlations) may not fully capture the models’ performance across all dimensions of language understanding. For example, the GMET task relies on manual evaluations by native speakers, which might introduce subjectivity and potential inconsistencies.
@@ A.3 WinoGrande (WG)
We translated & added the Winogrande dataset into our evaluation benchmark to assess models’ ability to perform commonsense reasoning and resolve linguistic ambiguities. The dataset’s adversarial construction reduces reliance on superficial statistical patterns, ensuring models rely on deep semantic and commonsense reasoning, which is essential for real-world applications like dialogue systems or conversational agents in Nepali. Including Winogrande in the Nepali, benchmark allows us to evaluate model strengths and limitations in handling nuanced linguistic structures and cultural contexts specific to Nepali, thereby improving their robustness for practical, context-sensitive applications.
@@ Appendix B Dataset Translation Approach
Given the unfunded nature of this research, we relied on personal resources and utilized the APIs of two large language models, GPT-4o-mini and Gemini2.5-flash, to translate datasets into Nepali. We processed data in batches of 50 to 100 rows, each containing text and its corresponding label, using automated scripts to manage batching, API interactions, and output collection. For tasks requiring nuanced understanding, such as Co-reference Resolution, manual translation and review were employed to ensure accuracy.
@@ B.1.1 Label Mismatch
Despite instructions to preserve labels, models produced correct Nepali translations that differed in meaning from the English source, resulting in label mismatches. In some cases, models also corrected errors in the original English, requiring manual review to ensure consistency. Examples are provided in Figure 15, Figure 16 and Figure 17.
@@ B.2 Suggestions on Translation
To improve translation quality and accuracy, we recommend the following strategies: [^2] Translation Repository
@@ E.1 Single Sentence Tasks
Best hyperparameter settings (model config, learning rate, epochs, batch size) for each model on SA (Sentiment Analysis), CoLA (Corpus of Linguistic Acceptability), and WG (Wino-Grande) tasks are reported in Table 6.

### v3

@@ Abstract
*Jinu Nyachhyon1,2, *Mridul Sharma1,2, Prajwal Thapa1,2, Bal Krishna Bal1 1Information and Language Processing Research Lab (ILPRL), Kathmandu University 2Institute for Research and Innovation in Intelligent Systems (IRIIS) Correspondence: nyachhyonjinu@gmail.com, mridulsharma3301@gmail.com
@@ 1 Introduction
SA 65.1K 16.3K Sentiment Macro F1, Acc Reviews, Tweets
@@
Nepali is written in the Devanagari script and is a highly inflected language. The Nepali language incorporates a complex system of noun, adjective, and verb inflections, including gender, case, and number Bal (2004). It has a rich vocabulary and is spoken in different dialects across various regions, and there are variations in vocabulary, grammar, and pronunciation. Developing and establishing robust models for Nepali requires reliable methods to evaluate their quality and effectiveness and it is essential to have tools that can assess how well these models address the language’s unique linguistic challenges while identifying their limitations.
@@ 2 Related Works
Nepali Sentiment Analysis (NepSA) Singh et al. (2020) is a targeted aspect-based sentiment analysis dataset, with 3,068 comments extracted from 37 YouTube videos across 9 channels. It is annotated using a binary sentiment polarity schema across six aspects: General, Profanity, Violence, Feedback, Sarcasm, and Out-of-scope. Another dataset, Aspect-Based Sentiment Analysis Tamrakar et al. (2020), contains 1,576 sentences, equally divided between positive and negative sentiments. Additional datasets, such as Nepali Language Sentiment Analysis - Movie Reviews Ghimire with 602 data points, and Nepali Sentiment Analysis Acharya with 2,161 data points found on Kaggle, are limited in size and domain-specific. For our benchmark, we utilized the NepCOV19Tweets dataset Sitaula et al. (2021), which includes ~33.5k sentiments labeled as positive, negative, or neutral. From these, we selected 14.9k positive and 13.5k negative data points for the SA dataset. A more recent dataset, Sentiment of Election-Based Nepali Tweets Pokharel , contains ~17.8k tweets but includes English characters and numbers, making it less suitable for our benchmarked dataset. To our knowledge, there are no publicly available datasets for coreference resolution, acceptability judgment, paraphrase and similarity detection, commonsense reasoning, pronoun coreference resolution, general masked evaluation, or NLI in the Nepali language. Despite some studies focusing on Nepali grammar, the lack of datasets for these advanced tasks limits the development of comprehensive NLU benchmarks.
@@ 8 Limitations
First, the datasets introduced in the NLUE benchmark were primarily created by translating existing English-language datasets from benchmarks such as GLUE and XGLUE, using automated tools like GPT-4o-mini and Gemini2.5-flash, supplemented by manual corrections. Although efforts were made to ensure translation accuracy, subtle linguistic nuances, cultural contexts, and idiomatic expressions specific to Nepali may not have been fully captured. The small size of certain datasets (e.g., CR and RTE) limits model performance and shows models’ lack of generalization on smaller datasets. Second, the study evaluates a range of models with varying parameter sizes (67M to 276M), but resource constraints prevented the inclusion of larger, state-of-the-art models or extensive hyperparameter tuning. Finally, the reliance on specific evaluation metrics (e.g., accuracy, Macro-F1 score, Spearman, and Pearson correlations) may not fully capture the models’ performance across all dimensions of language understanding. For example, the GMET task relies on manual evaluations by native speakers, which might introduce subjectivity and potential inconsistencies.
@@ A.3 WinoGrande (WG)
We translated & added the WinoGrande dataset into our evaluation benchmark to assess models’ ability to perform commonsense reasoning and resolve linguistic ambiguities. The dataset’s adversarial construction reduces reliance on superficial statistical patterns, ensuring models rely on deep semantic and commonsense reasoning, which is essential for real-world applications like dialogue systems or conversational agents in Nepali. Including WinoGrande in the Nepali benchmark allows us to evaluate model strengths and limitations in handling nuanced linguistic structures and cultural contexts specific to Nepali, thereby improving their robustness for practical, context-sensitive applications.
@@ Appendix B Dataset Translation Approach
Given the unfunded nature of this research, we relied on personal resources and utilized the APIs of two large language models, GPT-4o-mini and Gemini2.5-flash, to translate datasets into Nepali [^2] Translation Repository. We processed data in batches of 50 to 100 rows, each containing text and its corresponding label, using automated scripts to manage batching, API interactions, and output collection. For tasks requiring nuanced understanding, such as Co-reference Resolution, manual translation and review were employed to ensure accuracy.
@@ B.1.1 Label Mismatch
Despite instructions to preserve labels, models produced correct Nepali translations that differed in meaning from the English source, resulting in label mismatches. In some cases, models also corrected errors in the original English texts, requiring manual review to ensure consistency. Examples are provided in Figure 15, Figure 16 and Figure 17.
@@ B.2 Suggestions on Translation
To improve translation quality and accuracy, we recommend the following strategies:
@@ E.1 Single Sentence Tasks
Best hyperparameter settings (model config, learning rate, epochs, batch size) for each model on SA (Sentiment Analysis), CoLA (Corpus of Linguistic Acceptability), and WG (WinoGrande) tasks are reported in Table 6.